### v1

# Let and let be an open subset of Let and let be a family of smooth vector fields on which span the tangent space at every point of Let be positive integers. Let be the usual (Euclidean) norm on 5.1. Metric associated to a family of vector fields Let be an -dimensional vector space, and equip with the family of dilations defined by

Let and let be an open subset of Let and let be a family of smooth vector fields on which span the tangent space at every point of Let be positive integers. Let be the usual (Euclidean) norm on
5.1. Metric associated to a family of vector fields
Let be an -dimensional vector space, and equip with the family of dilations defined by
That is,
Fix a norm on which is homogeneous with respect to i.e. for all
Let Note that
We will write for the ball in with respect to the homogeneous norm
Since the Lebesgue measure of is proportional to
Let denote the linear map from to the Lie algebra of vector fields on given by
For any there exists such that the map defined by
exists and is a submersion. Here, is the exponential of vector fields (i.e., the flow at time ). We can see that exists for sufficiently small because is smooth. It is also clear that the derivative of at is surjective since spans the tangent space at Hence, exists and is a submersion for sufficiently small For further details, see [Goo76, Page 39].
Since is a submersion, the pre-image of a point is a (possibly empty) -dimensional submanifold of Note that whether or not is empty will depend on the choice of
Let be the Jacobian matrix of since is a submersion, is a -matrix of full rank. Define
Here, is the matrix transpose.
The co-area formula asserts that
Here is the -dimensional Hausdorff measure on the submanifold
See [Fed69, Theorem 3.2.12]. We will frequently refer to this measure and so we introduce the following notation.
Definition 5.1.
Let Let denote the following measure on
With this notation, the coarea formula for asserts that
We record a few essential features of
Lemma 5.2.
Let We have
and on
Proof.
By definition, if and only if and Since we have and hence if we have
This proves that if and then
By symmetry, we conclude the converse, and this proves the first claim.
The fact that on follows from (5.1).
β
Using we can also define the following function on
Definition 5.3.
Let If define
Otherwise,
Equivalently, if and if there exists such that Observe that is precisely the function denoted in [NSW85, Section 4]. The same metric was used by Street, see [Str23, Section 1.3].
We want to use as a quasi-metric on As defined, may fail to be a quasi-metric for the obvious reason that may be infinite. A more subtle problem is that is not necessarily symmetric: it might be that but
We deal with this issue by assuming that is as small as necessary. Theorem 2 of [NSW85] implies the following:
Theorem 5.4.
Every point has a neighbourhood such that for all and Moreover, for all
Henceforth, replacing with if necessary, we will make the following assumption:
Assumption 1.
Assume that is small enough so that is finite and symmetric on and also so that
Replacing by we assume without loss of generality that is constant for
We also assume that has compact closure, in order that all of the derivatives of with respect to are uniformly bounded over
Ultimately there will be no harm in making this assumption, since our final estimates will take place over compact subsets of which can be covered by a finite number of sets satisfying Assumption 1.
Using this new assumption, is a non-empty submanifold of for every
Now that is guaranteed to be finite and symmetric on we can use the following theorem.
Theorem 5.5 (Nagel-Stein-Wainger).
[NSW85, Section 4]
As given in Definition 5.3, defines a quasi-metric on
which is uniformly comparable to the Euclidean metric in the sense that there exist constants such that
If denotes the Lebesgue measure on then is a doubling quasi-metric measure space.
We will need some estimates on integrals with respect to the measures
Lemma 5.6.
Let and
There exists a constant such that
and also
If then
Proof.
If then (5.2), implies that so
It follows that for small the two claims in (1) are equivalent. The same follows for general by increasing the constant
By the coarea formula (5.1), if then
By increasing if necessary, this still holds for
This proves (1).
For (2), note that by definition if we have and so
β
5.2. Graded basis
Let be as above. In particular, we continue to enforce Assumption 1.
Let be an open subset of which is invariant under the -action defined by
The -tuple is said to a graded basis for the filtration defined by in the terminology of [AMY22, Definition 2.1].
Definition 5.7.
Let denote the projection map
A function on is said to be -properly supported if is proper on the support of That is, if for any compact subset of is compact.
Definition 5.9.
Let be -properly supported.
Assume that there exists such that
Assume that if and then is supported in
For fixed and define
That is, is a function on such that for any we have
Let
Our goal is to show that in Definition 5.9 obeys all of the conditions of Definition 3.1, with where is the Lebesgue measure. Conditions (I), (II) and (IV) are easy consequences of the Lemma 5.6, while Condition III will be more challenging.
Note that the condition that there exists such that
is equivalent to
Indeed, since is smooth there exists an -properly supported smooth function such that
and as
Lemma 5.10.
Let and be as in Definition 5.9. There exists a constant such that if
then That is, satisfies Condition (I) of Definition 3.1 with constant
Proof.
By assumption, we have
Hence, the result follows from Lemma 5.6.(2).
β
Lemma 5.11.
Let and be as in Definition 5.9. There exists a constant , independent of such that
That is, satisfies Condition (II) of Definition 3.1, with constant
Proof.
Exactly as in Lemma 5.10, we have
The result follows from Lemma 5.6.(1).
β
Lemma 5.12.
Let and be as in Definition 5.9. There exists a constant independent of such that
That is, satisfies Condition (IV) of Definition 3.1, with constant
Proof.
By the coarea formula (5.1) and the assumption on we have
β
5.3. Level sets of submersions
The final and most difficult condition to verify from Definition 3.1 is Condition (III). In order to achieve this, we need to understand how the measure varies with This can be done using classical vector calculus.
To simplify the notation, in this subsection we consider a general submersion
where recall that and and is equipped with a norm The fact that is equipped with a family of anisotropic dilations will not be important in this subsection. Ultimately we will consider which is only defined on an open subset of but for brevity we work here with defined on the entire space. Since our reasoning is local this is no loss of generality.
Given let denote the measure on given by
The coarea formula asserts that
In this subsection we will derive a formula for the difference
The intent will be to analyse the difference where is as in Definition 5.9.
The idea is to define a flow in which transports to
Let denote the vector field on given by
Since is a submersion and the vector field is nowhere vanishing. Let be the flow of this vector field. That is, or more specifically
solves the differential equation
In principle, the flow may only exist for a time
If it exists, the map transports to One way to see this is to define
The chain rule gives
Hence, if then for all
It follows that defines a smooth map from to In particular, at we have a smooth map
Similar reasoning shows that is a smooth inverse for so is a diffeomorphism.
The next step is to determine how changes with Again let We have
Here, is the pullback of to and is the determinant.
To see (5.5), let and By (5.3), we have
By a change of variables
Applying (5.3) again gives
Since and are arbitrary, (5.5) follows.
Using (5.5), we have
Therefore
Recall that where is the vector field (5.4). The Liouville-Ostrogradsky formula says that Jacobian determinant of obeys the differential equation
Since we have and hence
Thus,
With this preparation, we can prove the main result of this subsection.
Proposition 5.13.
Let be a submersion, let
There exists and such that for all ,
and we have
Proof.
Let be the vector field in (5.4).
Since is smooth and depends linearly on for every and and there is a constant such that
Let By the Picard-LindelΓΆf theorem, if
then for the flow exists and moreover
This implies that
Similarly,
Since
we have
Since this implies
Therefore,
Similarly,
Let Note that is supported in since
Now we apply (5.3), resulting in
By (5.8), (5.7)
and the numerical inequality
we have
Since we conclude the result with constant
β